# Generalized Guarantees for Variational Inference in the Presence of Even and Elliptical Symmetry

### v1

@@ 5 Discussion
Appendix A CONNECTION BETWEEN $\varphi$-DIVERGENCES AND $f$-DIVERGENCES
In this appendix, we review $f$-divergences and discuss their connection to $\varphi$-divergences. As in the rest of the paper, we assume $p$ and $q$ admit a density with respect to a Lebesgue measure.
Definition 14.
($f$-divergence Rényi, (1961)) We refer to a divergence as an $f$-divergence, if it can be written as
$D_{f}(p||q)=\int f\left(\frac{p(z)}{q(z)}\right)q(z)\text{d}z,$ (20)
where (i) $f(t)$ is convex, (ii) $|f(t)|<\infty$ for all $t>0$, (iii) $f(1)=0$, and $f(0)=\lim_{t\to 0^{+}}f(t)$.
The $f$-divergence is not a generalization of the $\varphi$-divergence, nor is it a special case thereof. Rather, each definition imposes different restrictions on a function that acts on the ratio $p(z)/q(z)$ inside the integral. There are many examples of divergences which are both $f$- and $\varphi$-divergences, including the divergences in Table 1, with one notable exception. The Rényi divergence of order $\alpha$,
$\text{D}_{\alpha}(p||q)=\int\frac{1}{\alpha(1-\alpha)}\left(\frac{p^{\alpha}(z)}{q^{\alpha}(z)}-1\right)\text{d}z,$ (21)
[…]
Another reason for working with the $\varphi$-divergences is because the regularity conditions in Section 2.4 concern the log-concavity of $p$. Hence it is more convenient to directly work with a function of $\log p$.
Appendix B SUPPORTING PROOFS
@@
B.1 Strict convexity of $D_{\varphi}(p||q_{\nu})$ in $\nu$
In this section, we show that if $p$ is somewhere-strictly log concave and $\varphi$ is monotone decreasing, then $D_{\varphi}(p||q_{\nu})$ is strictly convex in $\nu$. We begin with a weaker result.
@@
Building on Lemma 15, we obtain a result about the strict convexity of $D_{\varphi}(p||q)$.
@@
Assume the regularity conditions on $p$ and $q$ described in Section 2.4, and in particular that $p$ is log concave over $\mathbb{R}^{d}$. Consider the $\varphi$-divergence $D_{\varphi}(p||q)$ and suppose one of the following: (i) $p$ is somewhere-strictly log concave and $\varphi$ is strictly monotone decreasing, OR (ii) $\varphi$ is montone decreasing and strictly convex. Then $D_{\varphi}(p||q_{\nu})$ is strictly convex in $\nu$.
@@
Consider the function $g_{\zeta}:\mathbb{R}^{d}\to\mathbb{R}$, with $g_{\zeta}(\nu)=\log p(\zeta+\nu)-\log q_{0}(\zeta)$. We will now derive an inequality on $g_{\zeta}$, using Proposition 11 of Margossian and Saul, (2025). The latter only applies to functions with a univariate input and so we introduce $f_{\zeta}:\mathcal{R}\to\mathbb{R}$, a function whose domain $\mathcal{R}$ is the line that goes through $\nu_{0}$ and $\nu_{1}$. For $\nu\in\mathcal{R}$, $f_{\zeta}(\nu)=g_{\zeta}(\nu)$. Then, for $\zeta\in\Omega_{\lambda}$ and applying Proposition 11 in Margossian and Saul, (2025), we have
$f_{\zeta}(\nu_{\lambda})>(1-\lambda)f_{\zeta}(\nu_{0})+\lambda f_{\zeta}(\nu_{1}).$ (28)
@@
B.2 Proof for Guarantees in the Presence of Elliptical Symmetry
@@
The proof proceeds in two steps. First, we show that $D_{\varphi}(p||q_{\nu,S})$ is strictly convex in $S^{1/2}$ and therefore admits a unique minimizer $S$. Second, we derive a stationary point which has the desired form $S^{1/2}=\gamma M^{\frac{1}{2}}$ for some $\gamma>0$.
@@
$=$ $\int\varphi\left(\log\frac{p(S^{\frac{1}{2}}\zeta+\nu)}{|S|^{-\frac{1}{2}}q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta$
@@
$=$ $\int\varphi\left(\log p_{0}(M^{-\frac{1}{2}}(S^{\frac{1}{2}}\zeta)-\log q_{0}(\zeta)+\log|S^{\frac{1}{2}}|-\log|M^{\frac{1}{2}}|\right).$
@@
We now find a stationary point of $D_{\varphi}(p||q)$. Let $J=M^{-\frac{1}{2}}S^{\frac{1}{2}}$. Since $M^{-1/2}$ is fixed and invertible, the uniqueness of a solution for $S^{1/2}$ implies there is a unique solution for $J$. To complete the proof, we must show that $J=\gamma I$—with $I$ the identity matrix—is a stationary point for some $\gamma\in\mathbb{R}$. Since $q_{0}$ is spherically symmetric, we can define $g:\mathbb{R}^{+}\to\mathbb{R}$ such that
@@
To complete the proof, we show that the R.H.S of eq. (44) is continuous in $\gamma$. This follows from the assumption that $f^{\prime}$ and $\varphi^{\prime}$ are continuous. Therefore, there exists at least one value of $\gamma\geq 0$ such that eq. (44) is verified. But since $D_{\varphi}(p||q)$ has a unique minimizer with respect to $J$, $\gamma$ must be unique. Moreover, $J=\gamma I$ and so $S^{1/2}=\gamma M^{1/2}$, with $\gamma\geq 0$. ∎
B.3 Proof for Guarantees in the Presence of Partial Even Symmetry
@@
$=\int\frac{\nabla_{\zeta_{\sigma}}p(\nu+\zeta)}{q_{0}(\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta$
$=\int_{\bar{\sigma}}\int_{\sigma}\frac{\nabla_{\zeta_{\sigma}}p(\nu+\zeta)}{q_{0}(\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta_{\sigma}\text{d}\zeta_{\bar{\sigma}}.$ (45)
@@
we have that $p(z)$ is somewhere-strictly log concave in $z_{\sigma}$ over $\mathbb{R}^{|\sigma|}$ and applying Lemma 16 from Appendix B.1, we have that $D_{\varphi}(p||q_{\nu})$ is strictly convex in $\nu_{\sigma}$. Hence, any stationary point $\nu^{*}$ of $D_{\varphi}(p||q_{\nu})$ must verify $\nu^{*}_{\sigma}=\mu_{\sigma}$. ∎
B.4 Proofs for Guarantees in the Presence of Partial Elliptical Symmetry
@@
Let $\mathcal{Q}$ be the family of Gaussians. Suppose $p$ is elliptically symmetric along $\sigma$ about a point of even symmetry, which is linear in $z_{\bar{\sigma}}$, and with a constant normalized covariance matrix, and suppose $f$ is continuously differentiable and $|f^{\prime}(0)|<\infty$. In addition, suppose $p(z_{\sigma}|z_{\bar{\sigma}})$ is somewhere-strictly log concave over $\mathbb{R}^{|\sigma|}$. Then any minimizer of $\text{KL}(q||p)$ must verify $\nu_{\sigma|\bar{\sigma}}=\mu_{\sigma|\bar{\sigma}}$ and $S_{\sigma|\bar{\sigma}}=\gamma^{2}\Sigma_{\sigma|\bar{\sigma}}$ for some unique $\gamma\in\mathbb{R}$.
@@
B.5 Analysis of the Elliptical Funnel
@@
Appendix C EXPERIMENTAL DETAILS
@@
The code to reproduce all experimental results and figures in the paper is provided in the Supplemental Material. We use R as a scripting language and Stan (Carpenter et al.,, 2017; Stan Development Team,, 2025) as a probabilistic programming language to specify models and run VI and MCMC. Our work with Stan is greatly facilitated by the packages BridgeStan (Roualdes et al.,, 2023). All experiments are run on CPU using a 2.8 GHz Quad Core Intel Core i7 processor.
C.1 Targets
@@
This distribution is even and elliptically symmetric along its first two coordinates, $\sigma=x$. This is an extension of the two-dimensional Rosenbrock distribution by Roberts et al., (1997). The two-dimensional Rosenbrock distribution is even-symmetric along its first coordinate. We add an additional dimension in order to obtain a non-trivial partial elliptical symmetry.
schools $(d=10)$. A Bayesian hierarchical model of the effects of a preparation program for a standardized test across $N=8$ schools (Rubin,, 1981; Gelman et al.,, 2013). We observe $y_{i}$, the average change in test scores, and $\eta_{i}$, the empirical standard deviation across students, for each school. The model is then
@@
disease $(d=102)$. A model for the mortality counts across counties in Finland, due to alcoholism (Vanhatalo et al.,, 2010). For each county, we observe the mortality count, $y_{i}$, the standardized expected number of deaths, $y_{e,i}$, and the two-dimensional location of the country, $x_{i}$. The original model considers 911 counties, however we consider a random subset of 100 counties to reduce the computational cost of the experiment. The model uses a Gaussian process prior with a squared exponential kernel. Specifically, the prior covariance matrix $K$ is defined by,
@@
SKIM $(d=305)$. A sparse kernel interaction model (SKIM) (Agrawal et al.,, 2019). This model is a regularized regression model that accounts for interaction effect between covariates. Covariates are probabilistically selected using a horseshoe prior (Piironen and Vehtari,, 2017). As in Margossian et al., (2020), we apply the model to a genetic microarray classification data set on prostate cancer. We observe $N=102$ patients with $p=200$ pre-selected genetic covariates (out of a total of 5966 covariates) and denote $X\in\mathbb{R}^{N\times p}$ the design matrix. We observe for each patient $y_{i}$, a binary variable that indicates whether the patient has cancer.
@@
Following Agrawal et al., (2019), we marginalize out $\beta_{i}$ and $\beta_{ij}$ using a Gaussian process reparameterization. To define the Gaussian process’ covariance matrix $K$, we first introduce the matrices:
@@
C.2 VI algorithm
For the experiment in Section 4, we specify targets in Stan and fit them with ADVI (Kucukelbir et al.,, 2017). ADVI employs a Gaussian approximation over the unconstrained scale. Constrained variables are automatically transformed to an unconstrained scale by Stan. For example, a variable $z\in\mathbb{R}^{+}$ is replaced by $\log z\in\mathbb{R}$. In our experiments, we report estimates of the mean on the unconstrained scale, since our theoretical analysis applies to variables defined over $\mathbb{R}$. This choice allows us to test how predictive/illustrative our theory is, however practitioners may be more interested in estimates of the mean over the original scale.
@@ C.2 VI algorithm
C.3 Additional results on correlation

### v2

@@ Acknowledgment
Appendix A SUPPORTING PROOFS
@@
A.1 Strict convexity of $D_{f}(p||q_{\nu})$ for $\varphi$ convex and decreasing
In this section, we show that if $p$ is somewhere-strictly log concave and $\varphi(v)=f\circ\exp(v)$ is convex and monotone decreasing, then $D_{f}(p||q_{\nu})$ is strictly convex in $\nu$. We begin with a weaker result.
@@
Building on Lemma 14, we obtain a result about the strict convexity of $D_{f}(p||q)$.
@@
Assume the regularity conditions on $p$ and $q$ described in Section 2.4, and in particular that $p$ is log concave over $\mathbb{R}^{d}$. Consider the $f$-divergence $D_{f}(p||q)$ and suppose one of the following: (i) $p$ is somewhere-strictly log concave and $\varphi$ is strictly monotone decreasing, OR (ii) $\varphi(v)=f\circ\exp(v)$ is monotone decreasing and strictly convex. Then $D_{f}(p||q_{\nu})$ is strictly convex in $\nu$.
@@
Consider the function $g_{\zeta}:\mathbb{R}^{d}\to\mathbb{R}$, with $g_{\zeta}(\nu)=\log p(\zeta+\nu)-\log q_{0}(\zeta)$. We will now derive an inequality on $g_{\zeta}$, using Proposition 11 of Margossian:2025robust. The latter only applies to functions with a univariate input and so we introduce $h_{\zeta}:\mathcal{R}\to\mathbb{R}$, a function whose domain $\mathcal{R}$ is the line that goes through $\nu_{0}$ and $\nu_{1}$. For $\nu\in\mathcal{R}$, $h_{\zeta}(\nu)=g_{\zeta}(\nu)$. Then, for $\zeta\in\Omega_{\lambda}$ and applying Proposition 11 in Margossian:2025robust, we have
$h_{\zeta}(\nu_{\lambda})>(1-\lambda)h_{\zeta}(\nu_{0})+\lambda h_{\zeta}(\nu_{1}).$ (28)
@@
A.2 Proof for Guarantees in the Presence of Elliptical Symmetry
@@
The proof proceeds in two steps. First, we show that $D_{\varphi}(p||q_{\nu,S})$ is strictly convex in $S^{1/2}$ and therefore admits a unique minimizer $S$. Second, we derive a stationary point which has the desired form $S^{1/2}=\gamma M^{1/2}$ for some $\gamma>0$.
@@
$=$ $\int\varphi\left(\log\frac{p(S^{\frac{1}{2}}\zeta+\nu)}{q_{0}(\zeta)|S|^{-\frac{1}{2}}}\right)q_{0}(\zeta)\text{d}\zeta$
@@
$=$ $\int\varphi\left(\log p_{0}(M^{-\frac{1}{2}}S^{\frac{1}{2}}\zeta)-\log q_{0}(\zeta)+\log|S^{\frac{1}{2}}|-\log|M^{\frac{1}{2}}|\right).$
@@
We now find a stationary point of $D_{\varphi}(p||q)$. Let $J=M^{-\frac{1}{2}}S^{\frac{1}{2}}$. Since $M^{-1/2}$ is fixed and invertible, the uniqueness of a solution for $S^{1/2}$ implies there is a unique solution for $J$. To complete the proof, we must show that $J=\gamma I$—with $I$ the identity matrix—is a stationary point for some $\gamma\in\mathbb{R}$. Since $q_{0}$ is spherically symmetric, we can define $k:\mathbb{R}^{+}\to\mathbb{R}$ such that
@@
To complete the proof, we show that the R.H.S of eq. (44) is continuous in $\gamma$. This follows from the assumption that $g^{\prime}$ and $\varphi^{\prime}$ are continuous. Therefore, there exists at least one value of $\gamma\geq 0$ such that eq. (44) is verified. But since $D_{\varphi}(p||q)$ has a unique minimizer with respect to $J$, $\gamma$ must be unique. Moreover, $J=\gamma I$ and so $S^{1/2}=\gamma M^{1/2}$, with $\gamma\geq 0$. ∎
A.3 Proof for Guarantees in the Presence of Partial Even Symmetry
@@
$=\int\frac{\nabla_{\zeta_{\sigma}}p(\nu+\zeta)}{p(\nu+\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta$
$=\int_{\bar{\sigma}}\int_{\sigma}\frac{\nabla_{\zeta_{\sigma}}p(\nu+\zeta)}{p(\nu+\zeta)}\varphi^{\prime}\left(\log\frac{p(\nu+\zeta)}{q_{0}(\zeta)}\right)q_{0}(\zeta)\text{d}\zeta_{\sigma}\text{d}\zeta_{\bar{\sigma}}.$ (45)
@@
we have that $p(z)$ is somewhere-strictly log concave in $z_{\sigma}$ over $\mathbb{R}^{|\sigma|}$ and applying Lemma 15 from Appendix A.1, we have that $D_{\varphi}(p||q_{\nu})$ is strictly convex in $\nu_{\sigma}$. Hence, any stationary point $\nu^{*}$ of $D_{\varphi}(p||q_{\nu})$ must verify $\nu^{*}_{\sigma}=\mu_{\sigma}$. ∎
A.4 Proofs for Guarantees in the Presence of Partial Elliptical Symmetry
@@
Let $\mathcal{Q}$ be the family of Gaussians. Suppose $p$ is elliptically symmetric along $\sigma$ about a point of even symmetry, which is linear in $z_{\bar{\sigma}}$, and with a constant normalized covariance matrix, and suppose $g$ is continuously differentiable and $|g^{\prime}(0)|<\infty$. In addition, suppose $p(z_{\sigma}|z_{\bar{\sigma}})$ is somewhere-strictly log concave over $\mathbb{R}^{|\sigma|}$. Then any minimizer of $\text{KL}(q||p)$ must verify $\nu_{\sigma|\bar{\sigma}}=\mu_{\sigma|\bar{\sigma}}$ and $S_{\sigma|\bar{\sigma}}=\gamma^{2}\Sigma_{\sigma|\bar{\sigma}}$ for some unique $\gamma\in\mathbb{R}$.
@@
A.5 Analysis of the Elliptical Funnel
@@
Appendix B EXPERIMENTAL DETAILS
@@
The code to reproduce all experimental results and figures in the paper is provided in the Supplemental Material. We use R as a scripting language and Stan (Carpenter:2017; Stan:2025) as a probabilistic programming language to specify models and run VI and MCMC. Our work with Stan is greatly facilitated by the packages BridgeStan (Roualdes:2023). All experiments are run on CPU using a 2.8 GHz Quad Core Intel Core i7 processor.
B.1 Targets
@@
This distribution is even and elliptically symmetric along its first two coordinates, $\sigma=x$. This is an extension of the two-dimensional Rosenbrock distribution by Roberts:1997. The two-dimensional Rosenbrock distribution is even-symmetric along its first coordinate. We add an additional dimension in order to obtain a non-trivial partial elliptical symmetry.
schools $(d=10)$. A Bayesian hierarchical model of the effects of a preparation program for a standardized test across $N=8$ schools (Rubin:1981; Gelman:2013). We observe $y_{i}$, the average change in test scores, and $\eta_{i}$, the empirical standard deviation across students, for each school. The model is then
@@
disease $(d=102)$. A model for the mortality counts across counties in Finland, due to alcoholism (Vanhatalo:2019). For each county, we observe the mortality count, $y_{i}$, the standardized expected number of deaths, $y_{e,i}$, and the two-dimensional location of the country, $x_{i}$. The original model considers 911 counties, however we consider a random subset of 100 counties to reduce the computational cost of the experiment. The model uses a Gaussian process prior with a squared exponential kernel. Specifically, the prior covariance matrix $K$ is defined by,
@@
SKIM $(d=305)$. A sparse kernel interaction model (SKIM) (Agrawal:2019). This model is a regularized regression model that accounts for interaction effect between covariates. Covariates are probabilistically selected using a horseshoe prior (Piironen:2017). As in Margossian:2020, we apply the model to a genetic microarray classification data set on prostate cancer. We observe $N=102$ patients with $p=200$ pre-selected genetic covariates (out of a total of 5966 covariates) and denote $X\in\mathbb{R}^{N\times p}$ the design matrix. We observe for each patient $y_{i}$, a binary variable that indicates whether the patient has cancer.
@@
Following Agrawal:2019, we marginalize out $\beta_{i}$ and $\beta_{ij}$ using a Gaussian process reparameterization. To define the Gaussian process’ covariance matrix $K$, we first introduce the matrices:
@@
B.2 VI algorithm
For the experiment in Section 4, we specify targets in Stan and fit them with ADVI (Kucukelbir:2017). ADVI employs a Gaussian approximation over the unconstrained scale. Constrained variables are automatically transformed to an unconstrained scale by Stan. For example, a variable $z\in\mathbb{R}^{+}$ is replaced by $\log z\in\mathbb{R}$. In our experiments, we report estimates of the mean on the unconstrained scale, since our theoretical analysis applies to variables defined over $\mathbb{R}$. This choice allows us to test how predictive/illustrative our theory is, however practitioners may be more interested in estimates of the mean over the original scale.
@@ B.2 VI algorithm
B.3 Additional results on correlation